# Next Visit Diagnosis Prediction via Medical Code-Centric Multimodal Contrastive EHR Modelling with Hierarchical Regularisation

### v1

@@ Abstract
Predicting next visit diagnosis using Electronic Health Records (EHR) is an essential task in healthcare, critical for devising proactive future plans for both healthcare providers and patients. Nonetheless, many preceding studies have not sufficiently addressed the heterogeneous and hierarchical characteristics inherent in EHR data, inevitably leading to sub-optimal performance. To this end, we propose NECHO, a novel medical code-centric multimodal contrastive EHR learning framework with hierarchical regularisation. First, we integrate multifaceted information encompassing medical codes, demographics, and clinical notes using a tailored network design and a pair of bimodal contrastive losses, all of which pivot around a medical code representation. We also regularise modality-specific encoders using a parental level information in medical ontology to learn hierarchical structure of EHR data. A series of experiments on MIMIC-III data demonstrates effectiveness of our approach.
@@ 1 Introduction
To address the aforementioned characteristics of EHR data, we present a novel framework for Next Visit Diagnosis Prediction via Medical Code-Centric Multimodal Contrastive EHR Modelling with Hierarchical Regularisation (NECHO). To the best of our knowledge, this framework is the first work designed in a medical code-centric fashion for diagnosis prediction. It tightly and seamlessly entangles three distinct modalities of medical codes, demographics, and clinical notes through a meticulously designed multimodal fusion network and two bimodal contrastive losses. Its goal is to boost representational power by positioning demographics and clinical notes as supplementary modalities. Furthermore, we harness an auxiliary loss to regularise each modality-specialised encoder based on the ancestral level of medical codes, thereby successfully injects more general information from the ICD-9 medical concept ontology to improve the performance. Therefore, the main contributions of our work are threefold as follows:
@@ 3.2.1 Modality-Specific Feature Extraction
Before introducing our novel fusion strategies, we first explain modality-specific encoders that extract features from each modality. We design them as simple as possible to highlight the efficacy of our proposed fusion strategies. Thus, our framework is modular, with the potential for performance enhancement if the encoders are switched to more representative ones.
@@ 3.2.2 Multimodal Fusion Network
Self-Attention Transformer To extract sequential feature representations effectively and boost dependencies from the above two cross-modal and medical code representations, a self-attention transformer (SA) is employed. It processes across the single-patient visits:
$\hat{y}^{C}$ $=\text{SA}^{C}(\hat{H}^{C}),$ (7)
$\hat{y}^{C\to H}$ $=\text{SA}^{C\to H}(Z^{C\to H}),$
$\hat{y}^{C\to W}$ $=\text{SA}^{C\to W}(Z^{C\to W}).$
@@
Multimodal Adaptation Gate Rather than performing a simple concatenation of the three distinct representations, we modify and adopt previous multimodal adaptation gate (MAG) Rahman et al. (2020); Yang and Wu (2021) in the medical code-centric manner. First, we calculate the trimodal gating value $g\in\mathbb{R}$ and the displacement vector H by concatenating ($\oplus$) meaningful representations in the previous stage as:
$g=\text{Linear}([\hat{y}_{t}^{C}\oplus\hat{y}_{t}^{C\to H}\oplus\hat{y}_{t}^{C\to W}]),$ (9)
$\text{H}=\text{Linear}(g[\hat{y}_{t}^{C\to H}\oplus\hat{y}_{t}^{C\to W}]).$ (10)
This modification maximises the influence of medical code representation during the fusion process. Then, a weighted summation is performed between the medical code representation $\hat{y}_{t}^{C}$ and the displacement vector H to derive the multimodal representation M:
@@
$\hat{y}_{t+1}=\text{Sigmoid}(\text{Linear}(\text{M})),\$ (12)
$\mathcal{L}_{\text{{ce}}}$ $=\frac{1}{N}\frac{1}{T}\sum_{t=1}^{T}-\left({y}_{t+1}^{\text{{T}}}\log\hat{{y}}_{t+1}+\right.$ (13)
@@
where cross-entropy loss $\mathcal{L}_{\text{{ce}}}$ is applied as the loss function. $y_{t+1}$ is a ground truth with elements $|\mathbb{C}|$, which takes a value of 1 if the $i$-th code exists in $V_{t+1}$, otherwise 0. It is finally averaged over the mini-batch $N$.
@@ 3.2.3 Bimodal Contrastive Losses
$\mathcal{L}^{(m_{1},m_{2})}_{\text{bi-con}}=\frac{1}{N}\sum_{i=1}^{N}(\alpha l_{i}^{(m_{1}\to m_{2})}+(1-\alpha)l_{i}^{(m_{2}\to m_{1})}).$ (16)
@@ 3.3 Hierarchical Regularisation
$\mathcal{L}^{m}_{\text{{hrchy}}}$ $=\frac{1}{N}\frac{1}{T}\sum_{t=1}^{T}-\left({o}_{t+1}^{\text{{T}}}\log\hat{{o}}^{m}_{t+1}+\right.$ (19)
@@ 3.4 Model Optimisation
The final objective function $\mathcal{L}_{\text{total}}$ is a weighted sum of three loss terms: the cross-entropy loss $\mathcal{L}_{\text{ce}}$ between ground truth diagnosis and prediction, the medical code-centric two bimodal contrastive losses $\mathcal{L}_{\text{cont}}$, and the three modality-specific direct ancestral level hierarchical losses $\mathcal{L}_{\text{hrchy}}$.
@@
where $\lambda_{\text{ce}},\lambda_{\text{cont}}$, and $\lambda_{\text{hrchy}}$ are parameters that balance the different loss terms. The parameters of the model are updated via stochastic gradient descent (SGD) technique with respect to the calculated loss.
@@
Demo Demo. Extractor (Ours) 17.96 29.58 47.13 58.94
@@
NECHO w/ocode ${}_{\text{w/o}\>\text{code centering}}$ (Ours) 28.10 42.13 59.32 70.01
@@
Table 1: Experimental Results on MIMIC-III Data for Next Visit Diagnosis Prediction. Code, Demo, and Note are short for Medical Codes, Demographics, Clinical Notes, respectively. Best results are in boldface. 10k and 512 indicates the number of words. Unless specified otherwise, 10k words are processed for multimodal models with with clinical notes.
@@ 4.1.2 Implementation Details
Also, following the previous work (Radford et al., 2021), the temperature $\tau$ and alpha $\alpha$ are 0.1 and 0.25 for the contrastive loss. The coefficients of loss terms, $\lambda_{\text{ce}}$, $\lambda_{\text{con}}$, and $\lambda_{\text{hrchy}}$ are set to 1, 1, and 0.1, respectively. Especially, the $\lambda_{\text{hrchy}}$ is set relatively small to weakly regularise each modality-specific encoder to learn the parental levels of ICD-9 codes, without overly constraining them. We provide the experimental results on the different $\lambda_{\text{hrchy}}$ to Appendix C.
@@ 4.2.1 Next Visit Diagnosis Prediction Results
Furthermore, whilst NECHO does not completely surpass MIPO, replacing its simple medical code encoder with MIPO (NECHO w/ ${}_{\text{w/ MIPO}}$) outperforms MIPO. It especially achieves a 1.01% increase in top-30 accuracy, indicating that 1) our framework is modular, and 2) NECHO can predict additional accurate diseases than MIPO by leveraging complementary information from various modalities, emphasising its significance in real clinical settings. We provide a regarding case study to Section 4.2.3.
Another noteworthy point outside the multimodal strategies is that, amongst the clinical note baselines, Clinical BERT (Alsentzer et al., 2019) that is trained with a maximum of 512 tokens surpasses the combination model of BioWord2Vec (Zhang et al., 2019) and 1D CNN (Kim, 2014) with equivalent number of tokens but is inferior to that model trained with 10k tokens. This suggests that enhancing performance is more about increasing the number of tokens to process than model complexity in EHR learning. This also justifies our preference for BioWord2Vec over Clinical BERT within the realm of Pretrained Language Models.
@@ 4.2.2 Ablation Studies
Finally, the effectiveness of our novel parental level hierarchical regularisation is investigated. Its omission (w/o $\mathcal{L}_{\text{hrchy}}$) affects adversely model’s accuracy across various top-$k$ accuracies. This suggests that enforcing the encoders for three distinct modalities, guided by the parental levels of medical codes using an ICD-9 hierarchy, is essential for enhancing performance as it prevents the possible transmission of erroneous information when combining representations from distinct data modalities and injects the general information to encourage effective and accurate training.
@@
Preceding Demo
Age: 67, Gender: Male, Admission Type: Emergency, Admission Location: Transfer from hospital …
Codes
D96, D109, D97, D131, D101, D49, D110, D53, D138, D257
Notes
… he was taken to the Operating Room where mitral valve replacement was performed … Discharge Diagnosis: mitral valve mass … He experienced some visual hallucinations … IMPRESSION: 1. Enlarging bilateral pleural effusions. 2. Enlarging cardiac silhouette suspicious for a pericardial effusion, echocardiographic confirmation is suggested.
Subsequent Codes
D238, D53, D130, D106, D101, D49, D2, D3, D2616, D96
MIPO
D101, D128, D53, D108, D95, D259, D106, D131, D98, D55
NECHO
D96, D98, D101, D53, D138, D238, D49, D106, D2616, D663
@@ 4.2.3 Case Study
To qualitatively evaluate the predictive performance between MIPO (Peng et al., 2021) and our NECHO, we present a case study (Table 3) using a patient whose medical history shows a progression from a mitral valve issue to complications after surgery and cardiac rhythm disturbances. In the study, codes are formatted according to the Clinical Classifications Software (CCS) [^1] https://hcup-us.ahrq.gov/toolssoftware/ccs/AppendixASingleDX.txt and are sequenced based on their priority, significantly influencing the reimbursement for treatment. We prefix them with "D" to make them appear akin to diagnosis codes.
@@
Additionally, our model predicts D238 and D2616 using multifaceted information of both demographics and notes. D238 should be predicted for two points: 1) the patient was initially hospitalised due to emergency health problem according to demographics, and 2) his notes states visual hallucinations, monitoring for pericardial and pleural effusions. The prediction of D2616 aligns with potential risks associated with mitral valve replacement. On the contrary, MIPO’s prediction of D259 (Residual codes; unclassified) and D131 (Respiratory failure; insufficiency; arrest (adult)), which is considered less informative and a simple repetition from previous patient visits. D2 (Septicemia) and D3 (Bacterial infection) are not explicitly mentioned in the patient’s history, thus extremely challenging to predict.
Apart from multimodal EHR learning, the content following the "Impression" in the preceding notes is only explicitly found in radiology reports. This indicates the importance of considering all available clinical note types in next visit diagnosis prediction task to acquire a thorough understanding of a patient’s information.
Hence, this demonstrates the necessity of the effective multimodal fusion strategy for its capability of capturing complementary and unique information in other modalities, verifying the efficacy of the NECHO.
@@ 5 Conclusion
Next visit diagnosis prediction is beneficial in AI-driven healthcare applications and has shown remarkable progress. However, the multifaceted and hierarchical properties of EHR data are beyond the consideration for the most of existing studies. To address these limitations, we introduce the novel multimodal EHR modelling framework, NECHO. It effectively aggregates three heterogeneous modalities through meticulously designed multimodal fusion network and the pair of two bimodal contrastive losses in a medical code-centric manner. It also uses parental level information of ICD-9 codes to each modality-specialised encoder to learn more general information. Experimental results including ablation studies and case study on MIMIC-III data highlight the NECHO’s efficacy and superiority.
@@ 6 Limitations
From a data perspective, the model’s predictions are heavily biased to the training data. This means there’s a potential risk that the model might underperform when encountering patterns that is nonexistent in the dataset or originating from the different healthcare settings. Additionally, from a model perspective, firstly, the framework’s applicability is confined and has not been extended to a variety of clinical event prediction tasks, such as mortality, re-admissions, and length of stay, where different modalities might take main status. Secondly, it operates under the assumption that all data modalities are readily and consistently available for every patient. However, this assumption is impractical in that the availability of data can be compromised due to device malfunctions or human errors.
@@
7 Acknowledgement
We highly appreciate anonymous reviewers for their valuable comments that helped us to enhance quality and completeness of this manuscript.
@@ A.2 Feature Extraction Module for Demographics
We capture the non-stationary nature of the aforementioned attributes across clinical records at the individual level. For example, variables such as age and insurance type may change over time. Thus, we employ a single embedding layer $\text{{E}}_{\text{H}}^{n}$ to $n$-th attribute at $t$-th patient record, $h_{t}^{n}$. The features from each embedding layer are then concatenated ($\oplus$) and fed into a single linear layer paired with a ReLU activation function. It can be represented as:
$\bar{h}_{t}=\text{{E}}_{\text{H}}^{1}(h^{1}_{t})\oplus\text{{E}}_{\text{H}}^{2}(h^{2}_{t})\oplus\cdots\oplus\text{{E}}_{\text{H}}^{n}(h^{n}_{t}),$ (24)

### v5

@@ Abstract
Predicting next visit diagnosis using Electronic Health Records (EHR) is an essential task in healthcare, critical for devising proactive future plans for both healthcare providers and patients. Nonetheless, many preceding studies have not sufficiently addressed the heterogeneous and hierarchical characteristics inherent in EHR data, inevitably leading to sub-optimal performance. To this end, we propose NECHO, a novel medical code-centric multimodal contrastive EHR learning framework with hierarchical regularisation. First, we integrate multifaceted information encompassing medical codes, demographics, and clinical notes using a tailored network design and a pair of bimodal contrastive losses, all of which pivot around a medical codes representation. We also regularise modality-specific encoders using a parental level information in medical ontology to learn hierarchical structure of EHR data. A series of experiments on MIMIC-III data demonstrates effectiveness of our approach.
@@ 1 Introduction
To address the aforementioned characteristics of EHR data, we present a novel framework for Next Visit Diagnosis Prediction via Medical Code-Centric Multimodal Contrastive EHR Modelling with Hierarchical Regularisation (NECHO). To the best of our knowledge, this framework is the first work designed in a medical code-centric fashion for diagnosis prediction. It tightly and seamlessly entangles three distinct modalities of medical codes, demographics, and clinical notes through a meticulously designed multimodal fusion network and two bimodal contrastive losses. Its goal is to boost representational power by positioning demographics and clinical notes as supplementary modalities. Furthermore, we harness an auxiliary loss to regularise each modality-specialised encoder based on the ancestral level of medical codes, thereby successfully injecting more general information from the ICD-9 medical ontology. Therefore, the main contributions of our work are threefold as follows:
@@ 3.2.1 Modality-Specific Feature Extraction
Before introducing our novel fusion strategies, we first explain modality-specific encoders that extract features from each modality. We design them as simple as possible to highlight the efficacy of our proposed fusion strategies. In other words, our framework is modular, with the potential for performance enhancement if the encoders are switched to more representative ones.
@@ 3.2.2 Multimodal Fusion Network
Self-Attention Transformer To extract sequential feature representations effectively and boost dependencies from the above two cross-modal and medical code representations, a self-attention transformer (SA) is employed. It processes across the patient’s single visits:
$\hat{y}^{C}_{t}$ $=\text{SA}^{C}(\hat{H}^{C}_{t}),$ (7)
$\hat{y}^{C\to H}_{t}$ $=\text{SA}^{C\to H}(Z^{C\to H}_{t}),$
$\hat{y}^{C\to W}_{t}$ $=\text{SA}^{C\to W}(Z^{C\to W}_{t}).$
@@
Multimodal Adaptation Gate Rather than performing a simple concatenation of the three distinct representations, we modify and adopt previous multimodal adaptation gate (MAG) Rahman et al. (2020); Yang and Wu (2021) in the medical code-centric manner. First, we calculate the trimodal gating value $g\in\mathbb{R}$ and the displacement vector H by concatenating meaningful representations in the previous stage as:
$g=\text{Linear}(\text{concat}(\hat{y}_{t}^{C};\hat{y}_{t}^{C\to H};\hat{y}_{t}^{C\to W})),$ (9)
$\text{H}=\text{Linear}(g(\text{concat}(\hat{y}_{t}^{C\to H};\hat{y}_{t}^{C\to W}))).$ (10)
This modification maximises the influence of medical code representation during the multimodal fusion process. Then, a weighted summation is performed between the medical code representation $\hat{y}_{t}^{C}$ and the displacement vector H to derive the multimodal representation M:
@@
$\hat{y}_{t+1}=\text{Sigmoid}(\text{Linear}(\text{M})),$ (12)
$\mathcal{L}_{\text{{ce}}}$ $=\frac{1}{T}\sum_{t=1}^{T}-\left({y}_{t+1}^{\text{{T}}}\log\hat{{y}}_{t+1}+\right.$ (13)
@@
where cross-entropy loss $\mathcal{L}_{\text{{ce}}}$ is applied as the loss function. $y_{t+1}$ is a ground truth with elements $|\mathbb{C}|$, which takes a value of 1 if the $i$-th code exists in $V_{t+1}$, otherwise 0.
@@ 3.2.3 Bimodal Contrastive Losses
$\mathcal{L}^{(m_{1},m_{2})}_{\text{bi-con}}=\frac{1}{N}\sum_{i=1}^{N}(\alpha l_{i}^{(m_{1}\to m_{2})}+$ (16)
$(1-\alpha)l_{i}^{(m_{2}\to m_{1})}).$
@@ 3.3 Hierarchical Regularisation
$\mathcal{L}^{m}_{\text{{hrchy}}}$ $=\frac{1}{T}\sum_{t=1}^{T}-\left({o}_{t+1}^{\text{{T}}}\log\hat{{o}}^{m}_{t+1}+\right.$ (19)
@@ 3.4 Model Optimisation
The final objective function $\mathcal{L}_{\text{total}}$ is a weighted sum of three loss terms: the cross-entropy loss $\mathcal{L}_{\text{ce}}$ between ground truth diagnosis and prediction, the medical code-centric two bimodal contrastive losses $\mathcal{L}_{\text{bi-con}}$, and the three modality-specific direct ancestral level hierarchical losses $\mathcal{L}_{\text{hrchy}}$. It is formulated as:
@@
where $\lambda_{\text{ce}},\lambda_{\text{bi-con}}$, and $\lambda_{\text{hrchy}}$ are parameters that balance the different loss terms. The parameters of the model are updated via stochastic gradient descent (SGD) technique with respect to the calculated loss.
@@
Demo Demo Extractor (Ours) 17.96 29.58 47.13 58.94
@@
NECHO w/ocode ${}_{\text{w/o}\>\text{code centring}}$ (Ours) 28.10 42.13 59.32 70.01
@@
Table 1: Experimental Results on MIMIC-III Data for Next Visit Diagnosis Prediction. Code, Demo, and Note are short for Medical Codes, Demographics, Clinical Notes, respectively. Best results are in boldface. 10k and 512 indicates the number of words. Unless specified otherwise, 10k words are processed for multimodal models with clinical notes.
@@ 4.1.2 Implementation Details
Also, following the previous work (Radford et al., 2021), the temperature $\tau$ and alpha $\alpha$ are 0.1 and 0.25 for the contrastive loss. The coefficients of loss terms, $\lambda_{\text{ce}}$, $\lambda_{\text{bi-con}}$, and $\lambda_{\text{hrchy}}$ are set to 1, 1, and 0.1, respectively. Especially, the $\lambda_{\text{hrchy}}$ is set relatively small to weakly regularise each modality-specific encoder to learn the parental levels of ICD-9 codes, without overly constraining them. We provide the experimental results on the different $\lambda_{\text{hrchy}}$ to Appendix C.
@@ 4.2.1 Next Visit Diagnosis Prediction Results
Furthermore, whilst NECHO does not completely surpass MIPO, replacing its simple medical code encoder with MIPO (NECHO w/ ${}_{\text{w/ MIPO}}$) outperforms MIPO. It especially achieves a 1.01% increase in top-30 accuracy, indicating that 1) our framework is modular, and 2) NECHO can predict additional accurate diseases than MIPO by leveraging complementary information from various modalities, emphasising its significance in real clinical settings. We provide a regarding case study (Section 4.2.3).
Another noteworthy point outside the multimodal strategies is that, amongst the clinical note baselines, Clinical BERT (Alsentzer et al., 2019) that is trained with a maximum of 512 tokens surpasses the combination model of BioWord2Vec (Zhang et al., 2019) and 1D CNN (Kim, 2014) with equivalent number of tokens but is inferior to that model trained with 10k tokens. This suggests that enhancing performance is more about processing a large number of tokens than increasing model complexity in EHR learning. This also justifies our preference for BioWord2Vec over Clinical BERT within the realm of Pretrained Language Models.
@@ 4.2.2 Ablation Studies
Finally, the effectiveness of our novel parental level hierarchical regularisation is investigated. Its omission (w/o $\mathcal{L}_{\text{hrchy}}$) affects adversely model’s accuracy across various top-$k$ accuracies. This suggests that enforcing the encoders for three distinct modalities, guided by the parental levels of medical codes using an ICD-9 hierarchy, is essential for enhancing performance as it injects the general information and thus prevents the possible transmission of erroneous information when combining representations from distinct data modalities, thereby encouraging effective and accurate training.
@@
Preceding Demo Age: 67, Gender: Male, Admission Type: Emergency, Admission Location: Transfer from hospital …
Codes D96, D109, D97, D131, D101, D49, D110, D53, D138, D257
Notes … he was taken to the Operating Room where mitral valve replacement was performed … Discharge Diagnosis: mitral valve mass … He experienced some visual hallucinations … IMPRESSION: 1. Enlarging bilateral pleural effusions. 2. Enlarging cardiac silhouette suspicious for a pericardial effusion, echocardiographic confirmation is suggested.
Subsequent Codes D238, D53, D130, D106, D101, D49, D2, D3, D2616, D96
MIPO D101, D128, D53, D108, D95, D259, D106, D131, D98, D55
NECHO D96, D98, D101, D53, D138, D238, D49, D106, D2616, D663
@@ 4.2.3 Case Study
To qualitatively evaluate the predictive performance between MIPO (Peng et al., 2021) and our NECHO, we present a case study (Table 3) using a patient whose medical history shows a progression from a mitral valve issue to complications after surgery and cardiac rhythm disturbances. In the study, codes are formatted according to the Clinical Classifications Software (CCS) and are sequenced based on their priority, significantly influencing the reimbursement for treatment. We prefix them with "D" to make them appear akin to diagnosis codes.
@@
Additionally, our model predicts D238 and D2616 using multifaceted information of both demographics and notes. D238 should be predicted for two points: 1) the patient was initially hospitalised due to emergency health problem according to demographics, and 2) his notes states visual hallucinations, monitoring for pericardial and pleural effusions. The prediction of D2616 aligns with potential risks associated with mitral valve replacement. On the contrary, MIPO’s prediction of D259 (Residual codes; unclassified) and D131 (Respiratory failure; insufficiency; arrest (adult)), which is considered less informative and a simple repetition from previous patient visits. D2 (Septicemia) and D3 (Bacterial infection) are not explicitly mentioned in the patient’s history, thus extremely challenging to predict. Hence, this demonstrates the necessity of the effective multimodal fusion strategy for its capability of capturing complementary and unique information in other modalities, verifying the effectiveness of the NECHO.
Apart from multimodal EHR learning, the content following the "Impression" in the preceding notes is only explicitly found in radiology reports. This indicates the importance of considering all available clinical note types to acquire a thorough understanding of a patient’s information. This contrasts with previous findings (Hsu et al., 2020; Husmann et al., 2022) suggesting that certain specific note types are representative in EHR learning.
@@ 5 Conclusion
Next visit diagnosis prediction is beneficial in AI-driven healthcare applications and has shown remarkable progress. However, the multifaceted and hierarchical properties of EHR data are beyond the consideration for the most of existing studies. To address these limitations, we introduce the novel multimodal EHR modelling framework, NECHO. It effectively aggregates representations from three heterogeneous modalities through meticulously designed multimodal fusion network and the pair of two bimodal contrastive losses in a medical code-centric manner. It also uses parental level information of ICD-9 codes to regularise each modality-specialised encoder to learn more general information. Experimental results including the ablation studies and case study on MIMIC-III data highlight the NECHO’s efficacy and superiority.
@@ 6 Limitations
From a data perspective, firstly, the model’s predictions are heavily biased to the training data. This means there’s a potential risk that the model might underperform when encountering patterns that is nonexistent in the dataset or originating from the different healthcare settings. Secondly, it operates under the assumption that all data modalities are readily and consistently available for every patient. However, this assumption is impractical in that the availability of data can be compromised due to device malfunctions or human errors. Additionally, from a model perspective, the framework’s applicability is confined and has not been extended to a variety of clinical event prediction tasks, such as mortality, re-admissions, and length of stay, where different modalities might take main status.
@@
Acknowledgement
We highly appreciate anonymous EACL reviewers and area chairs for their valuable comments that helped us to enhance quality and completeness of this manuscript.
@@ A.2 Feature Extraction Module for Demographics
We capture the non-stationary nature of the aforementioned attributes across clinical records at the individual level. For example, variables such as age and insurance type may change over time. Thus, we employ a single embedding layer $\text{{E}}_{\text{H}}^{n}$ to $n$-th attribute at $t$-th patient record, $h_{t}^{n}$. The features from each embedding layer are then concatenated and fed into a single linear layer paired with a ReLU activation function. It can be represented as:
$\bar{h}_{t}=\text{concat}(\text{{E}}_{\text{H}}^{1}(h^{1}_{t});\text{{E}}_{\text{H}}^{2}(h^{2}_{t});\cdots;\text{{E}}_{\text{H}}^{n}(h^{n}_{t})),$ (24)